The formula S=91t+164 models the average annual salary, S, of major league baseball players, in thousands of dollars, t years after 1984. What was the average annual salary in 1994?
Answer by xcentaur(357)

S=91t+164
t=1994-1984=10
S=
=91(10)+164
=910+164
=1074
=1074 thousand dollars


Hope this helps,
best of luck.
[NB:in case the answer when cross-checked is wrong,drop me a line at xcentaur AT hotmail DOT com.Thanks.]